Minutes of MP3 Meeting of 5 May 2010
(Present: Nuria, Zinur, Knud, Reiner, Andrea, Andrzej, Michele, Bernhard, Markus, Adriaan, Edward, Mike)
- Information from LHC/CERN Management
It was decided that MP3 will not anymore report weekly to the CCC Monday Meeting but at LMC (on Wednesday afternoon). Nuria will ask for more details on expected content and form for these presentation.
It was anyway decided to KEEP the preparation of the weekly status by MP3 (now collected by the person in shift on Sunday morning), to discuss eventual hot subject at the MP3 Meeting in order to finalize the presentation for the LMC meeting on the following afternoon. Eventually, the collecting could be done by the Tuesday-afternoon and Wednesday-morning person in shift.
- MARKUS: on Status of LHC by MP3-CCC activities (see slides). Discussed points:
1) Statistics of interventions.
2) RQTF/D Feedback problem :
now it seems solved
3) RCBH19.L3B2 & RCBV19.L3B1:
now solved, circuits recovered for operation.
4) QPS_OK check for 600 A:
Now systematically checked as for the main circuits. To minimize the effect of recurrent spikes, after two consecutive reading (polled every 10 min) showing QPS_OK= BAD the Injection permit is removed. NB. a large fraction of the 600A circuits still have the periodic automatic reset during which QPS_OK is lost.
5) Lost QPS communication on Main circuits in S.34 on Friday morning:
Problem was due to a connection fault in one of the racks removed during the technical stop. The cables being too short, this problem is not excluded in the future.
6) Trips of QPS at 6h10 on 28 April where QH of 11 IPQ were fired:
The trip was originated in a electrical distribution stop. The magnets are covered by UPS but still saw a large spike in voltage and trigger the QPS detector. A software interlock may be necessary to order to avoid that the EIC miss such kind of events and powers the circuits without requesting the check of the discharges by MP3-CCC if not in shift.
- BERNHARD: on MPT (Magnet Performance Tracking)
He presented the new system put in place in the JIRA environment used by BE/CO (tape issues in a browser from CERN). Please refer to slides (also put into “shared documents” in MPE-QPS Repository).
- NURIA: on IPQ Splice Measurements
The measurement of 6kA bus-bar segments van be done through the ElQA device already in the tunnel for 8L. Jean-Philippe and Sandrine tested it successfully using the memo from R. Mompo. It is reminded to whomever is in shift, to check with the EiC the possibility to launch one test whenever the machine is available (no beam, no access in the adjacent sectors). Afterwards, the ELQA team will analyze it off-line. The goal is to scan all the machine during the next months. Take into account that the test take ~ 1 hour. The displacement of the equipment inside the machine will be done during technical stops or whenever possible following a well defined procedure by the ElQA team.
- REINER: Warning on Undulators U_RES drifting:
Following a request from G. Arduini on the RU.R4 circuit it does not seem necessary to change the QPS protection in this circuit. The drift in the u-res signal is extremely small and it is enough t o reset the detector once in about three weeks. Besides, there is a warning when U_res reaches 150 mV (threshold at 200mV) so that the circuit is reset in the next days. NB. Time: ~3 h ramp-down + 3 h ramp-up.
- KNUD: on status of new cards for RQTF/D:
They should be available for 19th of May. Afterwards they need to be tested. A slot for installation should be planned with OP.
Michele and Nuria
There are minutes attached to this event.
Show them.